# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ Abstract
Earlier in my career, prevailing approaches in the emerging field of market design largely reflected the experiences and perspectives of leaders who were commissioned to design or reform various institutions. Since being commissioned for a similar task seemed unlikely for me as an aspiring design economist, I developed my own minimalist approach to market design. By leveraging the policy objectives of stakeholders, my approach creates a new institution from the existing one with minimal interference in its elements that compromise the objectives.
Minimalist market design initially evolved through my holistic research and policy efforts in school choice (1997–2005) and kidney exchange (2003–2007). Given its success in these areas, I systematically applied this approach in various, often unconventional, real-world settings. In recent years, my efforts in minimalist market design led to the deployment of a liver exchange system in Turkey, which became a global leader within two years; the 2021 reform of the U.S. Army’s branching system for assigning cadets to military specialties; and the adoption of reserve systems during the COVID-19 pandemic for vaccine allocation in 15 states and therapy allocation in 2 states. This same methodology also predicted the rescission of a 1995 Supreme Court judgment in India, which had led to countless litigations and interruptions in public recruitment for 25 years, as well as the mandates of its replacement.
In this monograph, I describe the philosophy, evolution, and successful applications of minimalist market design while also presenting the advancements in matching theory achieved through these applications. In doing so, I provide a paradigm for economists who wish to influence policy and shape institutions through their research.
@@
the loving memory of my soulmate Banu Bedestenci Sönmez,
who left her homeland behind so that I could pursue my policy aspirations as a design economist
@@ 1 Minimalist Market Design: A Proven Framework
Beginning with the first spectrum auction by the Federal Communications Commission (FCC) in 1994 (Milgrom, 2000), [^1] Key figures in the design of the FCC’s spectrum auction include auction theorists such as John McMillan, who served as a consultant for the FCC; Paul Milgrom and Robert Wilson, who were consultants for Pacific Telesis; and Preston McAfee, who consulted for AirTouch Communications (McMillan, 1994; McAfee and McMillan, 1996). over the past three decades, the field of market design has emerged, with researchers in auction theory and matching theory playing increasingly active roles in the design and redesign of economic and social institutions.
In the decade following the FCC spectrum auction, other prominent applications of this emerging field included the redesign of the entry-level labor market for medical doctors in the U.S. (Roth and Peranson, 1999), the redesign of the school choice mechanisms in Boston and New York City (Abdulkadiroğlu and Sönmez, 2003; Abdulkadiroğlu et al., 2005a, b), and the design of a kidney exchange clearinghouse in New England (Roth et al., 2004, 2005a, 2005b).
@@ 1.1 What is Minimalist Market Design?
At its core, minimalist market design introduces solutions to failures in economic, political, and social institutions—whether through mechanisms, governing rules, or even less structured ad hoc systems—by surgically removing design flaws while adhering as closely as possible to current policy goals and the existing setup.
To illustrate this framework, consider an institution established to achieve multiple objectives, composed of several units, each with a specific role and interacting in complex ways. When such an institution fails to fulfill its mission, the shortcomings may stem from individual units or the interfaces between them. How can a design economist effectively address these failures?
@@
This is exactly how a design economist organizes their efforts and operates within the minimalist market design framework. [^2] Throughout this monograph, the term design economist is used interchangeably with market designer to refer to researchers in the market design community who use formal methods to study or design institutions. They are not necessarily trained as economists; many, in fact, come from fields such as computer science, operations research, and electrical engineering.
They start by identifying the mission of the institution—the primary objectives of decision makers, system operators, or other stakeholders involved in the institution’s design. Understanding the institution’s historical evolution is often crucial for recognizing these objectives.
@@
To seize this opportunity, the economist identifies the root causes of the failures, pinpointing problematic units or interactions. The economist then targets these root causes for correction—and nothing else—much like a surgeon performing a minimally invasive procedure. This focused approach bestows the “minimalist” characteristic upon the paradigm, distinguishing it from more mainstream approaches in market design.
@@
In the late 1990s, as a junior design economist, I observed that emerging market design approaches—for example, the engineering approach to market design (Roth and Peranson, 1999; Roth, 2002)—predominantly, if not exclusively, reflected the experiences of leading economists commissioned to design or reform institutions. [^3] Despite significant advancements in the field, this trend still persists. The commissioning of such projects by decision makers indicates their readiness for design or reform, often driven by urgent needs or crises, which makes them particularly receptive to proposed changes.
Recognizing that similar opportunities were unlikely to come my way soon, I realized that these experiences—shaped by more receptive decision makers—might not be as informative for design economists like myself, who aim to influence policy as critical outsiders. Therefore, I adopted an alternative approach that incorporated a robust persuasion strategy, one that could be effective even when decision makers have vested interests in maintaining the status quo—think of it as integrating political economy constraints into market design.
@@
In his thought-provoking article “Is the Scientific Paper a Fraud?”, Peter Medawar, winner of the 1960 Nobel Prize for Physiology or Medicine, criticizes the format of scientific writing (Medawar, 1963). He contends that “the scientific paper may be a fraud because it misrepresents the processes of thought that accompany or give rise to the work that is described in the paper.” Revisiting Medawar’s criticism, Howitt and Wilson (2014) argue that even in hypothesis-driven research, “the trend toward more explicit framing of a hypothesis is often misleading, as hypotheses may be framed to explain a set of observations post hoc, suggesting a linear process that does not describe the actual discovery.”
Although Medawar’s concern focuses on the discovery process in scholarly inquiry, similar issues arise when attributing policy successes to specific scholarly research in applied economics, particularly in market design. Often, the policy impact attributed to market design scholarship reflects post hoc scholarly progress that follows practical outcomes, especially when the design team is commissioned for the project. In this monograph, I will remain mindful of Medawar’s concerns and present the thought processes underlying minimalist market design. To this end, I will focus on applications where policy impact follows the initial research findings, presenting research and policy as a holistic interplay that captures their interactions and mutual evolution.
@@
Inspired by Medawar’s concerns, I will discuss two types of policy relevance in this monograph: policy impact and external validity.
@@
If a team of design economists successfully informs policy—and potentially changes an institution—directly through their research, I refer to this as policy impact. Crucially, in my definition, the research must precede the real-world impact. If the research is conducted after the practical impact has already occurred, although it may still constitute policy impact for the researcher, I find it problematic to attribute the impact to the follow-up research. The reason is similar to why p-hacking is considered a questionable practice in empirical and experimental research.
Sometimes, practical relevance follows the research but occurs independently of it, initiated by policymakers or other stakeholders who discover and pursue similar solutions. In the minimalist market design framework—where a key motivation is uncovering the institution’s intended design—this scenario is particularly likely. Strictly speaking, I do not classify such outcomes as policy impact. Instead, if decision makers independently act in ways consistent with the minimalist market design previously outlined in scholarly work—arriving at similar conclusions and implementing changes—I refer to this as external validity. So, whenever I speak of external validity, “life imitates science.”
@@
While achieving direct policy impact may be considered by many as the more “glorifying” outcome, I personally believe that external validity is at least as important and serves a unique role. Achieving policy impact—though much less likely for an outsider critic than for a commissioned insider—can be influenced by factors unrelated to the merits of the proposed design, such as personal charisma, strong connections, an urgent need to avert a crisis, or sheer luck. In contrast, external validity serves as direct proof that the minimalist framework is working exactly as intended. It strongly confirms that you have indeed addressed the root causes of a flawed institution’s failures and formulated a desirable improvement.
@@ 1.3 Evolution of Minimalist Market Design
Phase 2 emerged from that initial setback and spanned from the late 1990s to the late 2000s. During this period, minimalist market design evolved through a series of successes and challenges. Notable breakthroughs included work in school choice (Abdulkadiroğlu and Sönmez, 2003) and kidney exchange (Roth et al., 2004, 2005b, 2006, 2007). These achievements greatly benefited from the use-inspired theories developed in Phase 1, even when the original target applications differed. I began deploying customized use-inspired theories as a means of persuasion, which proved increasingly effective.
Phase 3—currently ongoing—marks my deliberate and systematic application of minimalist market design to shape both my research and policy initiatives, which began in the early 2010s. During this phase, I have applied the framework across various areas—often successfully, though not always—in collaboration with diverse teams, including:
@@ 1.4 Aims, Theses and Outline of the Monograph
First, I aim to outline a coherent institutional design framework—the minimalist market design—which adopts a holistic approach to research and policy. Intended for policy-oriented scholars who aspire to influence policy from outside the system, this framework enables even outsider critics to build persuasive cases for change grounded in practical, use-inspired basic research. The thesis associated with this aim is that by adopting minimalist market design, these scholars can increase the likelihood that their proposals will be adopted by decision-makers and implemented in practice.
@@
Third, I aim to provide a detailed intellectual history of key innovations and policy breakthroughs in the theory and practice of matching markets over the past 25 years. By connecting these innovations and breakthroughs through a series of Discovery-Invention Cycles à la Narayanamurti and Odumosu (2016), I also highlight the role of minimalist market design in these achievements.
@@
Section 2 introduces the foundations and key features of minimalist market design. Section 3 examines the interaction between theory and policy within the idealized framework of minimalist market design presented in this monograph. It also discusses the linear model of innovation outlined by Bush (1945), which shaped U.S. innovation policy for five decades following World War II, alongside two alternative paradigms: Pasteur’s Quadrant, proposed by Stokes (1997) and influential since the late 1990s, and Discovery-Invention Cycles, introduced by Narayanamurti and Odumosu (2016). Section 4 offers advice on conducting research and engaging with policy using minimalist market design and includes a practical guide. Sections 5 through 12 present a series of matching market applications, including: residential assignment (Section 5), school choice (Section 6), kidney exchange (Section 7), cadet-branch matching (Section 8), priority design via reserve systems (Section 9), affirmative action in India (Section 10), emergency rationing of scarce medical resources (Section 11), and liver exchange (Section 12). Section 13 presents insights, contributions, and future directions in behavioral market design. Finally, Section 14 concludes the monograph.
@@ 2 Foundations and Features of Minimalist Market Design
I have been fortunate to actively participate in this exciting field since the late 1990s. Thanks to minimalist market design, my efforts have resonated not only with the academic community but also with decision makers and laypeople across various areas.
@@ 2.1 Navigating the Challenges of Policy Influence as an Outsider Critic
Let’s begin by exploring the motivation behind minimalist market design. Why did I, as an aspiring design economist, venture into uncharted territory and experiment with an untested approach, rather than follow the methods established by the pioneers of this emerging field?
@@
At that time, most market design literature and methodologies focused on commissioned design or reform projects, where the need for change was already recognized, and expert input was actively sought—often in response to an urgent crisis. As an aspiring design economist, I realized that such opportunities might not come my way for a long time, if ever, posing a significant challenge to my policy ambitions. Moreover, I believed these experiences were less informative for situations where decision-makers had yet to acknowledge the need for change and were likely to have vested interests in preserving the status quo.
@@
Therefore, I questioned whether lessons from commissioned market design projects—typically written for an academic audience—would be equally effective for critics like myself seeking to inform policy from outside the system. While they might be, I had my doubts. Consequently, I developed a framework tailored to the needs of an aspiring design economist offering unsolicited policy advice as an outsider, often challenging established institutions.

### v4

@@ Abstract
Minimalist market design is an economic design framework developed from the perspective of an outsider—one seeking to improve real institutions without a commission or official mandate. It offers a structured, “minimally invasive” method for reforming institutions from within: identify their mission as understood by stakeholders, diagnose the root causes of failure, and refine only those elements that compromise that mission. By fixing what is broken and leaving the rest intact, the framework respects the tacit knowledge embedded in long-standing institutions, minimizes unintended consequences, and secures legitimacy that facilitates adoption.
Such targeted interventions often call for novel, use-inspired theory tailored to the institutional context. In this way, minimalist market design advances both theory and practice through a reciprocal process fostering collaboration across disciplines and between academic research and real-world practice.
Tracing the framework’s evolution over twenty-five years of intertwined progress in theory and real-world implementation across a range of matching market applications—including housing allocation, school choice, living-donor organ exchange for kidney and liver, military branch assignment in the U.S. Army, the allocation of vaccines and therapies during the COVID-19 pandemic, and the allocation of public jobs and college seats under India’s reservation system—this monograph reveals a consistent “less is more” ethos, showing how restrained, precisely targeted reforms can yield substantial policy improvements while advancing fundamental knowledge.
Forthcoming in the Econometric Society Monographs Series
@@
the loving memory of my soulmate Banu Bedestenci Sönmez, who left her homeland behind
so that I could pursue my policy aspirations as an economic designer
@@ 1 Minimalist Market Design: A Proven Framework
Beginning with the first spectrum auction held by the Federal Communications Commission (FCC) in 1994 (Milgrom, 2000, 2004; Klemperer, 2004), [^1] Key contributors to the design of the FCC spectrum auction included John McMillan (consultant to the FCC), Paul Milgrom and Robert Wilson (consultants to Pacific Telesis), and Preston McAfee (consultant to AirTouch Communications) (McMillan, 1994; McAfee and McMillan, 1996). the field of market design has emerged over the past three decades, with researchers in auction theory and matching theory playing increasingly active roles in designing and reforming economic and social institutions.
In the decade following the FCC spectrum auction, other prominent applications of this emerging field included reforms of the entry-level labor market for medical doctors in the U.S. (Roth and Peranson, 1999), and of school choice mechanisms in Boston and New York City (Abdulkadiroğlu and Sönmez, 2003; Abdulkadiroğlu et al., 2005a, b), as well as the design of a kidney exchange clearinghouse in New England (Roth et al., 2004, 2005a, 2005b).
@@ 1.1 What is Minimalist Market Design?
At its core, minimalist market design introduces solutions to failures in economic, political, and social institutions—whether through mechanisms, governing rules, or even less structured ad hoc systems—by surgically removing design flaws while adhering as closely as possible to current policy goals and the existing setup; Figure 1 summarizes the workflow.
Figure 1: Minimalist market design workflow: identify the mission, assess mission alignment, diagnose failures, target root causes, and implement a minimalist redesign—a scalpel, not a chainsaw.
To illustrate this framework, consider an institution established to achieve multiple objectives, composed of several units, each with a specific role and interacting in complex ways. When such an institution fails to fulfill its mission, the shortcomings may stem from individual units or the interfaces between them. How can an economic designer effectively address these failures?
@@
This is exactly how an economic designer organizes their efforts and operates within the minimalist market design framework. [^2] Throughout this monograph, the term economic designer is used interchangeably with market designer to refer to researchers in the market design community who use formal methods to study or design institutions. They are not necessarily trained as economists; many, in fact, come from fields such as computer science, operations research, and electrical engineering.
They start by identifying the mission of the institution—the primary objectives of decision-makers, system operators, or other stakeholders involved in the institution’s design. Understanding the institution’s historical evolution is often crucial for recognizing these objectives.
@@
To seize this opportunity, the economist identifies the root causes of the failures, pinpointing problematic units or interactions. The economist then targets these root causes for correction—and nothing else—much like a surgeon performing a minimally invasive procedure. [^3] This operational principle of minimalist market design resonates with Keynes’s famous remark: “If economists could manage to get themselves thought of as humble, competent people, on a level with dentists, that would be splendid!” (Keynes, 1930). This focused approach bestows the “minimalist” characteristic upon the paradigm, distinguishing it from more mainstream approaches in market design.
@@
In the late 1990s, as a junior economic designer, I observed that emerging market design approaches—for example, the engineering approach to market design (Roth and Peranson, 1999; Roth, 2002)—predominantly, if not exclusively, reflected the experiences of leading economists commissioned to design or reform institutions. The commissioning of such projects by decision-makers indicates their readiness for design or reform, often driven by urgent needs or crises, which makes them particularly receptive to proposed changes.
Recognizing that similar opportunities were unlikely to come my way soon, I realized that these experiences—shaped by more receptive decision-makers—might not be as informative for economic designers like myself, who aim to influence policy as critical outsiders. Therefore, I adopted an alternative approach that incorporated a robust persuasion strategy, one that could be effective even when decision-makers have vested interests in maintaining the status quo—think of it as integrating political economy constraints into market design.
@@
In his thought-provoking article “Is the Scientific Paper a Fraud?”, Peter Medawar, recipient of the 1960 Nobel Prize in Physiology or Medicine, criticized the conventional format of scientific writing (Medawar, 1963). He contended that “the scientific paper may be a fraud because it misrepresents the processes of thought that accompany or give rise to the work that is described in the paper.” This critique was later formalized by Kerr (1998) as HARKing—“Hypothesizing After the Results are Known.” [^5] The replication-crisis literature subsequently grouped HARKing with other questionable research practices, including p-hacking, selective reporting, and undisclosed flexibility in data collection (Simmons et al., 2011; John et al., 2012). In a similar vein, Howitt and Wilson (2014) argue that even in hypothesis-driven research, “the trend toward more explicit framing of a hypothesis is often misleading, as hypotheses may be framed to explain a set of observations post hoc, suggesting a linear process that does not describe the actual discovery.”
Although HARKing—as formalized by Kerr (1998) and anticipated by Medawar (1963)—was originally discussed in the context of discovery process in scholarly inquiry, similar challenges arise when attributing policy successes to specific scholarly research in applied economics, particularly in market design. Often, and I would argue misleadingly, the policy impact attributed to market design scholarship reflects post hoc scholarly progress that follows institutional changes, especially when the design team is commissioned for the project. [^6] The point is not to deny credit for the substantive contributions of commissioned designers, but to caution against attributing policy impact to the post hoc research they may later publish.
In this monograph, I aim to avoid such distortions in attributing policy impact to specific research by making explicit the underlying thought processes of minimalist market design. To this end, analogous to the use of pre-registration of research hypotheses in many fields, I focus on applications where policy impact both follows and aligns with the initial research findings, presenting research and policy as a holistic interplay that captures their interaction and mutual evolution.
@@
Accordingly, I discuss two types of policy relevance in this monograph: policy impact and external validity.
Policy Impact.
@@
If a team of economic designers successfully informs policy—and potentially changes an institution—directly through their research, often with their involvement but not necessarily always, I refer to this as policy impact. Crucially, in this definition, research must precede real-world impact. If the research is conducted instead after the practical impact has already occurred, akin to concerns about HARKing in empirical and experimental studies, I find it problematic to attribute the impact to the follow-up research. [^7] Impact may legitimately stem from a researcher’s earlier scholarship or practical involvement, but it should not be attributed to their subsequent post hoc work.
External Validity.
Sometimes, practical relevance comes after the research but occurs independently of it, when policymakers or other stakeholders develop and adopt solutions akin to those identified in earlier research. Strictly speaking, I do not classify such outcomes as policy impact; I refer to them instead as external validity. In this sense, whenever I speak of external validity, “life imitates science.”
In the minimalist market design framework—where a key motivation is to uncover the institution’s intended design and to recover it in the least disruptive way by addressing the root causes of failure—this scenario is particularly likely. If the minimalist exercise proves sound, decision-makers may independently reach similar conclusions and amend their institutions in ways that align with the underlying minimalist design.
@@
While achieving direct policy impact may be considered by many as the more “glorifying” outcome, I personally believe that external validity is at least as important and serves a unique role. Achieving policy impact—though much less likely for an outsider critic than for a commissioned insider—can be influenced by factors unrelated to the merits of the proposed design, such as personal charisma, strong connections, an urgent need to avert a crisis, or sheer luck. In contrast, external validity serves as direct proof that the minimalist framework is working exactly as intended. It strongly confirms that the analysis has addressed the root causes of an institution’s failures and formulated a desirable improvement.
@@ 1.3 Evolution of Minimalist Market Design
Phase 2 emerged from that initial setback and spanned from the late 1990s to the late 2000s. During this period, minimalist market design evolved through a series of challenges and successes. Notable breakthroughs included work in school choice (Abdulkadiroğlu and Sönmez, 2003) and kidney exchange (Roth et al., 2004, 2005b, 2006, 2007). These achievements greatly benefited from the use-inspired theories developed in Phase 1, even when the original target applications differed. I began deploying customized use-inspired theories as a means of persuasion, which proved increasingly effective.
Phase 3—which began in the early 2010s and is currently ongoing—marks my deliberate and systematic application of minimalist market design in shaping both my research and policy initiatives. In this phase, working with diverse teams, I have applied—often successfully, though not always—the framework across various areas, including:
@@ 1.4 Aims, Theses and Outline of the Monograph
First, I aim to outline a coherent economic design framework—the minimalist market design—which adopts a holistic approach to research and policy. Intended for policy-oriented scholars who aspire to influence policy from outside the system, this framework enables even outsider critics to build persuasive cases for change grounded in practical, use-inspired basic research. The thesis associated with this aim is that by adopting minimalist market design, these scholars can increase the likelihood that their proposals will be adopted by decision-makers and implemented in practice.
@@
Third, I aim to provide a detailed intellectual history of key innovations and policy breakthroughs in the theory and practice of matching markets over the past 25 years. By connecting these innovations and breakthroughs through a series of Discovery–Invention Cycles à la Narayanamurti and Odumosu (2016), I also highlight the role of minimalist market design in these achievements, as illustrated in Figures 36 and 37.
As a roadmap and preview, Table 1 compiles a broad set of minimalist designs discussed in this monograph that achieved policy relevance—both contributions from my research collaborations and notable designs by others—indicating the year of adoption, the form of policy relevance (whether direct policy impact, external validation, or commissioned design), and the triggers for adoption.
Table 1: Minimalist designs in the monograph that influenced real-world practice, with year of adoption, form of policy relevance of the underlying research, and adoption triggers.
Application Year Form of Policy Relevance Trigger of Adoption Section
Second-price Auction
  Stamp auctions 1893 Practitioner-driven (External Validation) Sealed bids; Parity with English auction 3.2.3
  New Zealand   Spectrum Auction 1989–90 Research-driven (Policy Impact) Transparency 3.2.3
FCC Spectrum Auctions (Collusion mitigation) 1997; 2008 Policy-maker-driven (External Validation) Tacit collusion; Transparency 14.1
School Choice
  Boston 2005 Research-driven (Policy Impact) Incentive compatibility 6.10
  England 2007 Policy-maker-driven (External Validation) Incentive compatibility 6.12.1
  Chicago 2009 Policy-maker-driven (External Validation) No justified envy; Incentive compatibility 6.12.2
Kidney Exchange
  NEPKE 2004 Research-driven (Policy Impact) Efficiency; Ethics 7.5
  3-way cycles 2006 Research-driven (Policy Impact) Efficiency 7.6
  NDD chains 2007 Research-driven (Policy Impact) Efficiency 7.6
AEA Job market (Signaling) 2006 Commissioned Coordination failures 14.2
Reputation Systems (Taobao) 2012 Research-driven (Policy Impact) Low response rate; Inflated ratings; Cold-start barrier 14.3
Platform Economy
  Airbnb 2016; 2018 Research-driven (Policy Impact) Racial bias 14.5
  Uber 2024 Research-driven (Policy Impact) Racial bias 14.5
Army Branching (USMA, ROTC) 2020 Research-driven (Policy Impact) No justified envy; Incentive compatibility; Transparency 8.9
Affirmative Action (India) 2020 Supreme Court mandate (External Validation) No justified envy 10.5
COVID-19 Resource Allocation 2020 Research-driven (Policy Impact) Ethics; Transparency 11.511.7
Liver Exchange (Turkey) 2022 Research-driven (Policy Impact) Efficiency; Ethics 12.6
@@
Section 2 introduces the foundations and key features of minimalist market design. Section 3 examines the interaction between theory and policy within the idealized framework of minimalist market design presented in this monograph. It also discusses the linear model of innovation outlined by Bush (1945), which shaped U.S. innovation policy for five decades following World War II, alongside two alternative paradigms: Pasteur’s Quadrant, proposed by Stokes (1997) and influential since the late 1990s, and Discovery–Invention Cycles, introduced by Narayanamurti and Odumosu (2016). Section 4 offers advice on conducting research and engaging with policy using minimalist market design and includes a practical guide. Sections 5 through 12 present a series of matching market applications, including: housing allocation (Section 5), school choice (Section 6), kidney exchange (Section 7), cadet–branch matching (Section 8), reserve systems (Section 9), affirmative action in India (Section 10), emergency rationing of scarce medical resources (Section 11), and liver exchange (Section 12). Section 13 presents insights, contributions, and future directions in behavioral market design. Section 14 highlights additional applications advanced by other researchers and practitioners across a wide variety of contexts, framed through the lens of minimalist market design. Finally, Section 15 concludes the monograph.
Throughout this monograph, I have integrated ideas and results from several of my earlier publications. The purpose here is not to simply re-present prior work, but to synthesize, extend, and reinterpret it under the overarching theme of minimalist market design. Where exposition or results closely parallel earlier work, the relevant articles are cited in the corresponding sections. In particular, several parts of the monograph draw directly on prior papers, including Section 5 on Abdulkadiroğlu and Sönmez (1999); Section 6 on Balinski and Sönmez (1999); Abdulkadiroğlu and Sönmez (2003); Pathak and Sönmez (2013); Section 7 on Roth et al. (2004); Sönmez and Ünver (2023); Section 8 on Sönmez and Switzer (2013); Greenberg et al. (2024); Section 9 on Dur et al. (2018b); Section 10 on Sönmez and Yenmez (2022, 2024); and Section 11 on Pathak et al. (2024).
@@ 2 Foundations and Features of Minimalist Market Design
I have been fortunate to actively participate in this exciting field since the late 1990s. Thanks to minimalist market design, my efforts have resonated not only with the academic community but also with decision-makers and laypeople across various areas.
@@ 2.1 Navigating the Challenges of Policy Influence as an Outsider Critic
Let’s begin by exploring the motivation behind minimalist market design. Why did I, as an aspiring economic designer, venture into uncharted territory and experiment with an untested approach, rather than follow the methods established by the pioneers of this emerging field?
@@
At that time, most market design literature and methodologies focused on commissioned design or reform projects, where the need for change was already recognized and expert input was actively sought—often in response to an urgent crisis. As an aspiring economic designer, I realized that such opportunities might not come my way for a long time if ever, posing a significant challenge to my ambition of making a positive impact through economic design. Moreover, I believed these experiences were less informative for situations where decision-makers had yet to acknowledge the need for change and were likely to have vested interests in preserving the status quo.
@@
Therefore, I questioned whether lessons from commissioned market design projects—typically written for an academic audience—would be equally effective for critics like myself seeking to inform policy from outside the system. While they might be, I had my doubts. Consequently, I developed a framework tailored to the needs of an aspiring economic designer offering unsolicited policy advice as an outsider, often challenging established institutions.